OpenStudy (anonymous):

If (a^x)^2/3 = 1/a^2, then x = (a) -1 (b) -2 (c) -3 (d) 2 (e) 1

OpenStudy (lgbasallote):

\[(a^x)^{2/3} = \frac{1}{a^2}\] \[\implies a^{2x/3} = a^{-2}\] therefore\[\frac{2x}3 = -2\] \[2x = -2 \times 3\] \[2x = -6\] \\[x = \frac{-6}2\] \[x = -3\] does that help?

OpenStudy (anonymous):

yesss thanks :)

OpenStudy (lgbasallote):

welcome